### v3

# Self-trapping phenomenon, multistability and chaos in open anisotropic Dicke dimer

Abstract
We investigate semiclassical dynamics of a coupled atom-photon interacting system described by a dimer of anisotropic Dicke model in the presence of photon loss, exhibiting a rich variety of non-linear dynamics. Based on symmetries and dynamical classification, we characterize and chart out various dynamical phases in a phase diagram. A key feature of this system is the multistability of different dynamical states, particularly the coexistence of various superradiant phases as well as limit cycles. Remarkably, this dimer system manifests self-trapping phenomena, resulting in a photon population imbalance between the cavities. Such a self-trapped state arises from a saddle-node bifurcation, which can be understood from an equivalent Landau-Ginzburg description. Additionally, we identify a unique class of oscillatory dynamics “self-trapped limit cycle”, hosting self-trapping of photons. The absence of stable dynamical phases leads to the onset of chaos, which is diagnosed using the saturation value of the decorrelator dynamics. Moreover, the self-trapped states can coexist with chaotic attractor, which may have intriguing consequences in quantum dynamics. Finally, we discuss the experimental relevance of our findings, which can be tested in cavity and circuit quantum electrodynamics setups.
Introduction: Nonequilibrium dynamics of the quantum many body systems has become a subject of intense research in recent years rev_noneq_1 ; rev_noneq_2 ; Moore_noneq_3 ; Huse_noneq_4 . The advancement of ultracold atomic system as well as cavity and circuit quantum electrodynamics (QED) paves the way for exploring a range of nonequilibrium phenomena, including various types of phase transitions and thermalization processes rev1 ; rev2 ; rev4 ; Serge_Haroche ; Carusotto_Review ; Helmut_Ritsch ; Esslinger_rev ; Zollar_rev ; Steven_Girvin_1 ; Houck ; Review_ergodicity_1 ; Blatter ; Le_Hur ; knap1 ; jcref7 ; fazio1 . These atom-photon interacting systems inherently exhibit dissipative effects due to various loss processes, providing a platform to study the rich dynamical behavior of open quantum systems Daley_dissipation ; Mueller_dissipation ; Hendric ; Keeling_Dicke ; Hemmarich ; Dis_tran4 ; Dis_tran5 ; Dis_tran6 ; Dis_tran7 ; Dicke_5 ; Dicke_6 ; Dis_tran9 ; Dis_tran10 ; Dis_tran11 . Understanding nonequilibrium phases of open quantum systems and nature of the transition between them has attracted significant interest Dis_tran1 ; Dis_tran8 ; Chitra ; Cuiti_dist_trans_1 ; Zollar_dist_trans ; dissipative_transition_CQED ; Cuiti_dist_trans_2 ; Milies_dist_trans . Seminal experiments have demonstrated the formation of exotic phases and nonequilibrium transitions by coupling cold atoms with cavity field supersolid_1 ; supersolid_2 ; Hemmarich ; A_M_Rey_cavity_nature ; A_M_Rey_cavity_exp ; Dissipative_Rydberg ; B_L_Lev ; ADM_1 ; Dicke_exp_1 ; Dicke_exp_2 ; Dicke_exp_3 ; Stojan_Rebic , as well as in the circuit QED array dissipative_transition_CQED . Additionally, oscillatory dynamics in the form of quantum limit cycles (LC) typically occur in the presence of dissipation, which leads to the intriguing time crystalline phase LC_Disssipative_Rydberg ; Limit_cycle_Hemmerich_2 ; continous_time_crystal ; Quantum_vanderpol ; quantum_LC_4 ; quantum_LC_5 ; quantum_LC_6 ; Dicke_correlator . On the other hand, the signature of chaos and its manifestation in the context of ergodicity and thermalization is a well studied area for isolated quantum system rev_noneq_2 ; Review_ergodicity_1 , which is however less explored for open quantum systems in the presence of dissipation Haake_chaos_1 ; Casati_chaos_1 ; Zyczkowsky ; Shepelyansky ; Prosen_3 ; BHT_chaos ; Deb_TCH ; Minganti ; Russomanno . Furthermore, the collective nature of these atom-photon systems allows to analyze the dynamics semiclassically, offering a pathway to probe the onset of chaos in open quantum systems Casati_chaos_1 ; QPT_Chaos_Brandes ; Sudip_review ; Graham_EPL ; Kolovsky .
The light-matter interacting system in a cavity containing multiple two level atoms can effectively be described by the anisotropic Dicke model ADM_1 ; ADM_2 ; Dis_tran4 . The paradigmatic Dicke model QPT_0 and its variants exhibit a plethora of fascinating phenomena, which include various types of phase transitions, chaos and thermalization QPT_Chaos_Brandes ; Haake_2012 ; Lea_2023 ; A_M_Rey_Dicke_1 ; Scar_Dicke ; Scar_sudip ; Dicke_1 ; Dicke_2 . In the present work, we consider nonequilibrium dynamics of two Josephson coupled cavities containing many atoms described by a dimer of anisotropic Dicke model. The photon hopping between the cavities can be engineered JCH_Plenio ; npj_hopping , as already demonstrated in the circuit QED experiments. In the presence of photon loss, this open anisotropic Dicke dimer (ADD) model exhibits a rich variety of dynamical phenomena as well as the formation of different nonequilibrium phases and transitions between them. In this letter, we focus on a new class of steady-state and oscillatory dynamics that lead to self-trapping phenomena of light even in the presence of photon loss, detectable from the photon number imbalance between the cavities. Self-trapping phenomena have attracted interest since their observation in cold atom systems Obarthalar_1 ; Oberthalar_2 ; AC_DC , and have been explored both theoretically and experimentally in other systems as well Self_trapping1 ; JC_dimer_expt ; Shenoy1 ; Shenoy2 ; Sebastian_ST ; JCD_Houck ; Vivek_JCD ; Manas_TCD . However, while such states typically have a finite lifetime in an open environment Self_trapping1 ; Shenoy_damping ; Scar_sudip , the present model (ADD) displays a stable self-trapping phenomena even in the presence of photon loss. Furthermore, we investigate the dissipative chaos in ADD and its physical consequences.
Model and semiclassical analysis: We consider two coupled cavities each containing $N$ two level atoms interacting with a single cavity modes (as shown in Fig.1(a)) which can be described by the Dicke dimer model with the following Hamiltonian,
$\hat{\mathcal{H}}\!\!$ $=$ $\!\!\!-J\left(\hat{a}_{\mathrm{L}}^{\dagger}\hat{a}_{\mathrm{R}}+\hat{a}_{\mathrm{R}}^{\dagger}\hat{a}_{\mathrm{L}}\right)\!+\!\!\sum_{i=\rm L,R}\left[\omega\hat{a}_{i}^{\dagger}\hat{a}_{i}+\omega_{0}\hat{S}_{zi}\right.$ (1)
$\!\!\!+\left.\frac{\lambda_{-}}{\sqrt{2S}}\left(\hat{a}_{i}^{\dagger}\hat{S}_{i-}+\hat{a}_{i}\hat{S}_{i+}\right)+\frac{\lambda_{+}}{\sqrt{2S}}\left(\hat{a}_{i}\hat{S}_{i-}\!+\!\hat{a}_{i}^{\dagger}\hat{S}_{i+}\right)\!\right]$
where the site index $i=\rm L(R)$ represents left(right) cavity, $\hat{a}_{i}$ annihilates photon mode with frequency $\omega$, and $J$ is the hopping amplitude of the photons between the cavities. Each cavity contains $N$ two level atoms with energy gap $\omega_{0}$, that are collectively represented by large spins $\hat{\vec{S_{i}}}$ with magnitude $S=N/2$, and $\lambda_{\pm}$ are atom-photon coupling strengths of the anisotropic Dicke model (ADM), which has been extensively studied both classically and quantum mechanically Dis_tran4 ; Dis_tran5 ; Dis_tran6 ; Dis_tran7 .
For this system with $S\gg 1$, the scaled operators $(\hat{x}_{i}+\imath\hat{p}_{i})/\sqrt{2}=\hat{a}_{i}/\sqrt{S}$ and $\hat{\vec{s}}_{i}=\hat{\vec{S}}_{i}/S$ become classical, as they satisfy, $[\hat{x}_{i},\hat{p}_{i}]=\imath/S$ and $[\hat{s}_{ai},\hat{s}_{bi}]=\imath\epsilon_{abc}\hat{s}_{ci}/S$, where $1/S$ plays the role of reduced Planck constant.
Realistically, photon loss from the cavities is inevitable, resulting in non-unitary time evolution of the density matrix (DM) of this open ADD model, which can be described by the Lindblad master equation Breuer ,
$\frac{d\hat{\rho}}{dt}=-\iota[\hat{H},\hat{\rho}]+\sum_{i=\rm L,R}\left(2\hat{L}_{i}\hat{\rho}\hat{L}_{i}^{\dagger}-\{\hat{L}_{i}^{\dagger}\hat{L}_{i},\hat{\rho}\}\right),$ (2)
where $\hat{L}_{i}=\sqrt{\kappa}\hat{a}_{i}$ account for the most dominant source of dissipation in cavity QED setups due to the photon decay with amplitude $\kappa$ ADM_1 ; Dicke_exp_1 ; Dicke_exp_2 ; Dicke_exp_3 ; Limit_cycle_Hemmerich_2 . In the absence of the coupling $J$, the above master equation has already been used to describe the single cavity experiments ADM_1 ; Dicke_exp_1 ; Dicke_exp_2 ; Dicke_exp_3 ; Limit_cycle_Hemmerich_2 . Apart from relaxation dynamics, the nonlinearity in the Lindblad master equation can also drive the open quantum system to a variety of fascinating stationary and non-stationary states Drossel . Throughout the paper, we set $\hbar,k_{B}=1$ and scale the coupling strengths (time) by $J$ $(1/J)$. The time evolution of the average value of any operator $\hat{O}$ can be obtained using the relation $\frac{d\langle\hat{O}\rangle}{dt}={\rm Tr}(\hat{O}\dot{\hat{\rho}})$. Within the mean-field approximation, the expectation of the product of operators can be decomposed as $\langle\hat{A}\hat{B}\rangle=\langle\hat{A}\rangle\langle\hat{B}\rangle$, which is valid for $S\rightarrow\infty$ Sudip_review ; Dis_tran8 ; Dis_tran11 ; Dicke_correlator .
Figure 1: (a) Schematic of the open anisotropic Dicke dimer. (b) Spin orientation representing different dynamical classes.
For the typical experimental setups with $N\approx 10^{4}$, the Eq.(2) can be analyzed semiclassically ADM_1 ; Dicke_exp_1 ; Dicke_exp_2 ; Dicke_exp_3 ; Limit_cycle_Hemmerich_2 and the dynamics of the scaled classical observables is described by following equations of motion (EOM),
[…]
where $\bar{i}\neq i$ and $\alpha_{i}=(x_{i}+\imath p_{i})/\sqrt{2}$ with scaled photon number $n_{i}=|\alpha_{i}|^{2}$. For convenience, we define the dynamical variables by an array $\mathcal{X}_{i}=\{Q_{i},s_{zi}\}$, with $Q_{i}=\{x_{i},p_{i},s_{xi},s_{yi}\}$. The above EOM obey $\mathbb{Z}_{2}$ parity symmetry corresponding to $\{Q_{\rm L},Q_{\rm R}\}\rightarrow\{-Q_{\rm L},-Q_{\rm R}\}$ and $s_{zi}\rightarrow s_{zi}$. Additionally, the EOM remain invariant when the site indices are interchanged, $\rm L\leftrightarrow\rm R$, exhibiting exchange symmetry. Both symmetries are relevant for characterizing different dynamical states.
To understand the overall dynamics and different nonequilibrium phases, we analyze the steady states (fixed points) $\mathcal{X}^{*}_{i}$ from the EOM and their stability, as discussed in the supplementary material SM . Introducing the dynamical variables, $\mathcal{X}_{\pm}=\{\frac{Q_{\rm L}\pm Q_{\rm R}}{2},\frac{s_{z{\rm L}}\pm s_{z{\rm R}}}{2}\}$ helps to classify the phase space dynamics in the following categories:
i) Symmetric: as evident from the EOM (see Eq.(3)), $\{Q_{-},s_{z-}\}=0$ remains as a fixed point, which can, in turn, impose a constraint and the dynamics is governed by the variables $\mathcal{X}_{+}$ only, within a restricted phase space.
ii) Anti-symmetric: on the other hand, the fixed point $\{Q_{+},s_{z-}\}=0$ confines the dynamics to a restricted subspace of the dynamical variables $\mathcal{X}_{-}$.
Note that $s_{z{\rm L}}=s_{z{\rm R}}$ for both classes, however, the projection of spins in the $x-y$ plane remains parallel (anti-parallel) for symmetric (anti-symmetric) class, as depicted in Fig.1(b) foontnote .
iii) Mixed: generally, the dynamics may not be restricted to the symmetric or anti-symmetric classes only, especially when the corresponding fixed points become unstable, resulting in their mixing.
Figure 2: Symmetric and anti-symmetric superradiant phases and limit cycles: (a) Classical phase diagram of stability regions of SSR (blue), ASR (pink) phases, their bistability (cyan) and the coexistence of ASR, SLC (violet) and SLC, ALC (lime), as well as the NP↑↑ (yellow stars) and NP↓↓ (blue diamonds). Transition of SSR (ASR) from NP↓↓ represented by pink (black) lines. Normal phases NP↑↓, NP↓↑ are always unstable. (b) Quantum dynamics within stochastic wavefunction approach for $S=5$: average scaled photon number $\langle\hat{n}_{\rm L}\rangle/S$ for different quantum trajectories converge to the classical steady state of either SSR or ASR in the bistable regime at $\lambda_{-}=2.2,\lambda_{+}=1.2$ (magenta star in (a)). The classical spin dynamics depicting the coexistence of (c) ASR and SLC at $\lambda_{-}=1.2,\lambda_{+}=1.64$ (green triangle) and (d) both limit cycles SLC, ALC for $\lambda_{-}=1.4,\lambda_{+}=2.0$ (violet circle). All energies (time) are measured by $J(1/J)$. We set $\hbar,k_{B}=1$ and $\omega=2.7,\omega_{0}=0.9,\kappa=0.9$ for all figures.
Next, we investigate the steady states characterized by the above mentioned dynamical classes and their stability. Interestingly, the EOM corresponding to the symmetric and anti-symmetric classes reduce to that of a single ADM with effective photon frequency $\omega_{\mp}=\omega\mp J$ respectively. However, their stability can be quite different as the fluctuations are not restricted to any particular symmetry class. First, we identify two stable normal phases (NP↓↓,NP↑↑) with zero photon number and the spin polarization $s^{*}_{z+}=\mp 1$ respectively. By changing the coupling strength, the normal phase NP↓↓ undergoes a transition to parity symmetry broken superradiant phase with non-zero photon number. In contrast to the single ADM, in this case, both the symmetric (SSR) and anti-symmetric (ASR) superradiant phases emerge from NP↓↓, each belonging to their respective dynamical class. The stability regimes of the normal phases and the formation of stable SR phases as well as their transitions in the parameter space are summarized in the phase diagram, given in Fig.2(a). The nature of the SR transition changes from a pitchfork to a saddle-node bifurcation depending on the parameter regime, as depicted by solid and dashed lines in Fig.2(a) respectively. This scenario of nonequilibrium transition can be captured analytically from an effective Landau-Ginzburg potential $\mathcal{F}(m)=\frac{a}{2}m^{2}+\frac{b}{4}m^{4}+\frac{c}{6}m^{6}$, with the order parameter $m^{2}=1+s_{z+}$ SM . For $b>0$, the continuous transition (pitchfork) changes to a first-order transition (saddle-node) at the tricritical point (TP) with $a,b=0$.
Interestingly, both the SSR and ASR phases can coexist in a large region of the phase diagram, exhibiting bistability (Fig.2(a)). However, their basins of attraction do not overlap, guiding the appropriate choice of initial conditions for the detection of the respective phases.
At the upper boundary of the bistable regime (dashed-dotted line) in the phase diagram in Fig.2(a)), the SSR phase becomes unstable and undergoes a Hopf bifurcation, giving rise to a pair of limit cycles belonging to the symmetric dynamical class (SLC), around two unstable branches of SSR phase. Further increasing $\lambda_{+}$, the ASR phase also becomes unstable and undergoes a similar Hopf bifurcation, giving rise to another pair of anti-symmetric LC (ALC). Interestingly, the SLC can coexist with either the ASR phase or the ALC in different parameter regions, giving rise to fascinating bistability. In these regions, trajectories converge to different attractors depending on initial conditions, as shown by the spin dynamics on the Bloch sphere (Fig.2(c,d)). Eventually, these limit cycles become unstable, leading to a quasi-periodic motion and chaos.
Figure 3: Self-trapping phenomenon: (a) Stability regions of steady states (see the text for details). The solid and dashed blue lines indicate the critical lines for the ST state. The stable ST state is bounded by the Dashed dotted black line. Dynamics of the photon imbalance $Z_{p}$ of (b1) ST state at $\lambda_{-}=1.15$, $\lambda_{+}=2.4$ (green square in (a)) and (b2) the STLC for $\lambda_{-}=1.15$, $\lambda_{+}=2.12$ (magenta circle in (a)). (c) Shape of the STLC (on the Bloch sphere) in the left (blue) and right (red) cavities at the parameter same as panel (b2). (d) Spin Dynamics reaching different steady states based on initial conditions in the multistability regime at $\lambda_{-}=0.8$, $\lambda_{+}=2.9$ (orange star in (a)).
Self-trapping phenomenon: To this end, we focus on another type of mixed class of dynamics, which breaks the exchange symmetry between two cavities, leading to a fascinating self-trapping phenomenon of photons. In the self-trapped (ST) state, the photon population in one of the cavities is larger, resulting in a non-vanishing photon number imbalance $Z_{p}=\frac{n_{\rm L}-n_{\rm R}}{n_{\rm L}+n_{\rm R}}$ and unequal spin polarization $|s^{*}_{z{\rm L}}|\neq|s^{*}_{z{\rm R}}|$ SM . The signature of the stable ST state can be detected from the dynamics of the population imbalance $Z_{p}$, which saturates to a non-vanishing value (see Fig.3(b1)). The nature of the transition leading to the ST state and its regime of stability are summarized in Fig.3(a). To investigate the formation of the ST steady state, we find that it stems from either of the unstable normal phases NP↑↓ or NP↓↑ with $|s^{*}_{z-}|=1$ and $s^{*}_{z+}=0$. Similar to the superradiant transition, the ST state originates from the aforementioned normal phases through a pitchfork bifurcation (continuous transition) along a critical line (solid blue line in Fig.3(a)), which eventually changes its nature to a saddle-node bifurcation (dashed blue line in Fig.3(a)), above a tricritical point. The phase boundary and the nature of transitions to the ST state can also be understood by constructing an effective LG potential $\mathcal{F}(m)$, where $m^{2}=1-s_{z-}$ serves as an order parameter (see SM for details). We emphasize that the stable ST phase originates only from the saddle-node bifurcation and exists in a specific region of the phase diagram, as shown in Fig.3(a). Another type of ST state originating from the unstable SR phase, always remains unstable, which is omitted from our discussion.
Interestingly, within a large region of the phase diagram shown in Fig.3(a), the stable ST state coexists with both the stable superradiant phases SSR and ASR, as well as with the stable normal phase NP↑↑, leading to an intriguing multistability of nonequilibrium phases. For such coexisting fixed points, the different phases can be probed dynamically depending on the initial conditions, as shown in Fig.3(d). The different fixed points in this multistable region have a non-overlapping basin of attraction SM , which guides the choice of the initial state suitably to observe the self-trapping phenomenon. Notably, the stable self-trapped state persists for a wide range of decay rates $\kappa$ SM , ensuring the robustness of the phenomenon even in the presence of dissipation.
Apart from the ST fixed point, we also identify a unique oscillatory dynamics leading to the self-trapping phenomenon, which we dub as self-trapped limit cycle (STLC). Such a limit cycle exists within a narrow region close to the boundary of the ST state (see Fig.3(a)), which is typically observed by initializing the dynamics close to the unstable normal phase NP↑↓. As seen from Fig.3(c), the spin dynamics corresponding to the self-trapped limit cycle leads to the formation of two distinctly different contours on the Bloch sphere of two cavities, exhibiting the exchange symmetry breaking. Furthermore, the self-trapping phenomenon corresponding to STLC is evident from the dynamics of photon imbalance $Z_{p}$, which shows periodic oscillation with non-vanishing average value (see Fig.3(b2)).
Figure 4: Dissipative Chaos: (a) Average Lyapunov exponent $\bar{\Lambda}$ on the phase diagram as a color scale. (b) Average Lyapunov exponent (left axis) and the saturation value of decorrelator $D_{s}$ (right axis) as a function of $\lambda_{+}$ for $\lambda_{-}=1.0$, exhibiting different dynamical regimes. Chaotic dynamics of (c) the spin in one cavity over Bloch sphere and (d) the photon imbalance $Z_{p}$ at the interaction strengths $\lambda_{-}=0.7,\lambda_{+}=1.4$.
Chaotic dynamics : In addition to the various steady states and limit cycles, the ADD model also exhibits dissipative chaos in a narrow region of the phase diagram (see Fig4(a)). Due to the instability of the fixed points, the region of regular dynamics shrinks significantly, which leads to the onset of chaos. The chaos stems from the instability of the symmetric limit cycles (SLC), leading to the chaotic motion of spins revolving around two pairs of unstable SSR fixed points (see Fig.4(c)), resembling the behavior of the Lorentz system Lorentz ; Strogatz .
We identify the chaotic region (depicted in Fig.4(a)) from the positive average Lyapunov exponent (LE) $\bar{\Lambda}$, computed for an ensemble of phase space points Lichtenberg ; Lyapunov . Furthermore, we study the decorrelator Decorr1 ; Decorr2 ; Decorr3 ; Decorr4 ; Decorr5 ; Decorr6 ; Decorr7 dynamics of spin $D(i,t)=1-\langle{\bf s}_{i}^{a}\cdot{\bf s}_{i}^{b}\rangle$, where the initial configurations $a,b$ differ by a small perturbation, $\mathbf{s}^{b}=\mathbf{s}^{a}+\delta\mathbf{s}$, and the averaging is performed over the initial states. In the chaotic regime, the decorrelator initially grows and finally saturates below unity, whereas it remains vanishingly small for regular phase space dynamics. As evident from Fig.4(b) the saturation value of the decorrelator $D_{s}$ serves as an alternate measure to quantify overall chaoticity, which can efficiently identify the chaotic region in the parameter space. Fig.4(a) depicts the gradual transition of the chaotic region into quasi-periodic motion, characterized by a peaked distribution of frequencies and a small value of the Lyapunov Exponent (LE).
Figure 5: Coexistence of chaos and self-trapped state: (a) Trajectories on the Bloch sphere converging to the two branches of ST state (black) corresponding to the parity symmetry and chaotic attractor (red). Dynamics of $s_{z-}$ for (b1) of ST state and (b2) chaotic regime. We set $\lambda_{-}=1.0$ and $\lambda_{+}=2.25$ from the region marked in Fig.4(b).
In the current scenario of dissipative chaos, the attractor occupies a finite region of phase space. In consequence, the ST steady states can coexist with chaos, as shown over the Bloch sphere in Fig.5(a). This remarkable coexistence phenomena can be probed dynamically, as the trajectory initialized with zero photon number and $s_{z{\rm L}}=1,s_{x{\rm R}}=\pm 1$ converges to the ST state, while chaotic dynamics can be observed for initial $N_{\downarrow\downarrow}$ state (compared in Fig.5(b1,b2)). Note that the degree of chaos and self-trapping phenomenon can also depend on the detuning $\delta=\omega-\omega_{0}$ SM .
Quantum dynamics: To this end, we study the semiclassical dynamics using truncated Wigner approximation (TWA) considering a large number of atoms, where a stochastic noise is added due to photon loss continous_time_crystal ; Limit_cycle_Hemmerich_2 ; quantum_LC_6 ; TWA_1 ; TWA_2 ; A_M_Rey_TWA . We demonstrate the signature of bistability between SSR and ASR phases within TWA, where the dynamics is attracted to either of the SR phases depending on the initial conditions. Furthermore, the self-trapping phenomenon of photons is also verified from the TWA simulation SM . Interestingly, the self-trapped limit cycle also persists in presence of quantum fluctuations up to a certain time scale, which increases with increasing number of atoms. The onset of chaos can also be unveiled from the broad frequency spectrum of individual stochastic trajectories.
Moreover, we confirm these intriguing phenomena from the full quantum dynamics, within the stochastic wavefunction approach Q_traj_1 ; Q_traj_2 for a small number of atoms. In the bistable regime of SSR and ASR phases, the quantum trajectories converge to either of the classical steady states (as shown in Fig.2(b)). Even for small number of atoms, we observe that a fraction of quantum trajectories are attracted towards the self-trapped state, confirming its signature, which we discuss in SM .
Discussion: The anisotropic Dicke dimer with photon loss displays diverse dynamical phenomena, such as the multistability of nonequilibrium phases, limit cycles, onset of chaos, and quasi-periodic motion. Notably, we have identified steady states as well as a new type of limit cycle leading to the self-trapping phenomenon, offering the possibility of observing a unique photonic time crystal. The saturation value of the decorrelator can serve as an alternate measure for chaoticity. The impact of quantum fluctuations on the dynamical phases deserves further investigation, which may induce a dynamical transition between the coexisting phases. The presence of a chaotic attractor may facilitate chaos-assisted tunnelling between the self-trapped states, opening avenues for exploring the quantum signature of dissipative chaos.
In conclusion, this atom-photon dimer system offers a platform to study a wide range of intriguing dynamical phenomena, including multistability, photon self-trapping, and dissipative chaos, relevant to experimental setups in cavity and circuit QED with potential applications in quantum technology.
Acknowledgements.
Acknowledgments: We thank Sudip Sinha for comments and fruitful discussions. D.M. acknowledges support from Prime Minister Research Fellowship (PMRF).